### v2

@@ 4.1 The generalization of the replica method
If $\omega(T,X)$ takes different constant values at the two endpoints, denoted as $\omega_{a}$ and $\omega_{-a}$, respectively. The contributions to the integral (52) from different endpoints correspond to spheres with radii $\omega_{a}X_{a}$ and $\omega_{-a}X_{a}$. Therefore, these integrals should be replaced by the partition function on spheres with radii $\omega_{a}X_{a}$ and $\omega_{-a}X_{a}$, respectively.
@@ 4.2 System with finite size
In the previous subsection, we reviewed the CHM mapping method for calculating the entanglement entropy of the $T\bar{T}$-deformed CFT in spatial intervals on a two-dimensional plane. Specifically, we computed the entanglement entropy for a subsystem on the boundary of the dS spacetime in Poincaré coordinates. We now extend this method to two-dimensional cylindrical and spherical surfaces to examine whether the $T\bar{T}$ deformation introduces corrections to the entanglement entropy for finite-size systems. The corresponding holographic models for these geometries are AdS and dS spacetimes in global coordinates, respectively. The entanglement entropy calculation for general spatial intervals in curved space has been discussed in ref. Grieninger:2019zts .
@@ 4.2.1 AdS case
$S_{A}=$ $\frac{c_{\rm AdS}}{3}\mathrm{arctanh}\,\left(\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r_{\rm eff}^{2}}}}\right)$
@@
After the $T\bar{T}$ deformation, its holographic boundary is a 2-dimensional sphere determined by $t=t_{c}$. The subregion $A$ is defined by the arc $\Theta=\pi/2$, $\Phi\in(-\Phi_{a},\Phi_{a})$ on this 2-dimensional sphere. As done previously, we begin by mapping the sphere to the complex plane using the transformation
$\Theta=2\arctan\sqrt{T^{2}+X^{2}}\,,\quad\Phi=\frac{1}{2i}\ln\frac{T+iX}{T-iX}\,.$ (80)
@@
$ds^{2}$ $=\frac{\ell_{\rm dS}^{2}\cosh^{2}t_{c}\sin^{2}\Phi_{a}}{(1+\cos(\phi-\Phi_{a})\cos(\Phi_{a})\sin(\theta))^{2}}(d\theta^{2}+\sin^{2}\theta d\phi^{2})$
@@
$\Omega(\theta,\phi)=\frac{1}{1+\cos(\phi-\Phi_{a})\cos(\Phi_{a})\sin(\theta)}\,.$ (84)
@@
$S_{A}$ $=-\frac{ic_{\rm dS}}{3}\mathrm{arccoth}\,\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\rm eff}^{2}}}}\right)+\frac{\pi c_{\rm dS}}{6}$
@@
This result is consistent with the area of the RT surface calculated in Appendix A.2, and also matches the findings of ref. Chen:2023eic when $\Phi_{a}=\frac{\pi}{2}$.
@@ 4.3 System with finite temperature
$S_{A}=\frac{c_{\rm AdS}}{6}\left[\mathrm{arctanh}\,\left(\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r_{{\rm eff},\theta=0}^{2}}}}\right)+\mathrm{arctanh}\,\left(\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r_{{\rm eff},\theta=\pi}^{2}}}}\right)\right]\,.$ (96)
@@
where we have replaced $R_{0}$ with $\frac{2\pi\ell_{\rm AdS}^{2}}{\beta}$. In the high-temperature limit $\beta\ll 1$, the last term is negligible compared to the second term. This result is consistent with the findings in refs. Chen:2018eqk ; Apolo:2023ckr by identifying $R_{c}^{2}=\frac{6\ell_{\rm AdS}^{2}}{\mu\pi c_{\rm AdS}}$.
5 (A)dS/dS case and non-locality
In this section, we apply the methods from the previous section to the AdS/dS, dS/dS, and half-dS holography cases. For all three, the holographic boundary is a timelike hypersurface. In dS/dS and half-dS holography, previous studies Kawamoto:2023nki ; Geng:2019ruz suggest that the dual field theory may exhibit non-locality. Unlike the examples discussed earlier, the boundary-induced metrics in this section are time-dependent. As a result, the properties of entanglement entropy may vary significantly over time, potentially influencing the determination of the strong subadditivity or the boosted strong subadditivity. This issue will be discussed in detail in this section.
5.1 AdS/dS
For the AdS/dS correspondence, the bulk metric is expressed as
$ds^{2}=\ell_{\rm AdS}^{2}(dw^{2}+\sinh^{2}w(-dt^{2}+\cosh^{2}tdx^{2})),$ (98)
where $x\sim x+2\pi$, and the holographic boundary is determined by $w=w_{c}\to\infty$. We choose the boundary subregion $A$ as $x\in(-x_{a},x_{a})$ at $t=t_{0}$. In the Euclidean signature, after rescaling the time coordinate, the boundary metric can be written as
$ds^{2}=\ell_{\rm AdS}^{2}\frac{\sinh^{2}w_{c}}{\cosh^{2}\tilde{t}}\left(d\tilde{t}^{2}+dx^{2}\right)\,,$ (99)
where $\tilde{t}$ is related to $t$ by the relation
$\tilde{t}=\ln\frac{\cot\frac{t}{2}+1}{\cot\frac{t}{2}-1}\,.$ (100)
This metric is conformally equivalent to the cylinder. As we have done in the last section, we can map it to the complex plane through the following transformations
and the resulting conformal flat metric is
$ds^{2}=\ell_{\rm AdS}^{2}\sinh^{2}w_{c}\frac{4\left(dT^{2}+dX^{2}\right)}{\left(1+T^{2}+X^{2}\right)^{2}}\,.$ (102)
The two endpoints of the subregion $A$ in the $(T,X)$ coordinates are
$(T_{1},X_{1})=(e^{\tilde{t}_{0}}\cos x_{a},-e^{\tilde{t}_{0}}\sin x_{a}),\quad(T_{2},X_{2})=(e^{\tilde{t}_{0}}\cos x_{a},e^{\tilde{t}_{0}}\sin x_{a})\,.$ (103)
Using the CHM map,
$T\rightarrow\frac{e^{\tilde{t}_{0}}\sin x_{a}\sin\theta\sin\phi}{1+\sin\theta\cos\phi}+e^{\tilde{t}_{0}}\cos x_{a}\,,\quad X\rightarrow\frac{e^{\tilde{t}_{0}}\sin x_{a}\cos\theta}{1+\sin\theta\cos\phi}\,,$ (104)
$ds^{2}=\frac{4e^{2\tilde{t}_{0}}\ell_{\rm AdS}^{2}\sinh^{2}w_{c}\sin^{2}x_{a}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)}{\left(1+\sin\theta\cos\phi+e^{2\tilde{t}_{0}}\sin\theta\cos(\phi-2x_{a})+e^{2\tilde{t}_{0}}\right)^{2}}\,.$ (105)
It is easy to verify that the conformal factor at the endpoints $\theta=0,\pi$ is the same constant, which can be absorbed into $r_{\rm eff}$. Therefore, the effective radius is
$r_{\text{eff}}=\ell_{\rm AdS}\sinh w_{c}\cos t_{0}\sin x_{a}\,,$ (106)
or, in Lorentz signature, $r_{\text{eff}}=\ell_{\rm AdS}\sinh w_{c}\cosh t_{0}\sin x_{a}$. Finally, the entanglement entropy of the subsystem is
$S_{A}\simeq\frac{c_{\rm AdS}}{3}\log(2\sinh w_{c}\cosh t_{0}\sin x_{a})+\frac{c_{\rm AdS}}{12\sinh^{2}w_{c}\cosh^{2}t_{0}\sin^{2}x_{a}}\,.$ (107)
If we choose the UV cutoff as $\sinh w_{c}=1/\epsilon$, we find that eq.(107) agrees with the result in ref. Kawamoto:2023nki . Alternatively, if we select the subregion as $t=t_{0}$, $x\in[-\frac{\pi}{2},\frac{\pi}{2}]$, the entanglement entropy is $S\simeq\frac{c_{\rm AdS}}{3}w_{c}$, which matches the result in ref. Donnelly:2018bef . For verification, we holographically calculate the area of the RT surface, which is expressed as
$D_{12}$ $=\ell_{\rm AdS}\mathrm{arccosh}(1+2\sinh^{2}w_{c}\cosh^{2}t_{0}\sin^{2}x)$
$\simeq 2\ell_{\rm AdS}\log(2\sinh w_{c}\cosh t_{0}\sin x_{a})+\frac{\ell_{\rm AdS}}{2\sinh^{2}w_{c}\cosh^{2}t_{0}\sin^{2}x_{a}}.$ (108)
After dividing $D_{12}$ by $4G$ and substituting the relation $G=3\ell_{\rm AdS}/2c_{\rm AdS}$, it is consistent with (107) up to the second-order correction of the $T\bar{T}$ deformation.
5.2 dS/dS
@@ 5.1 AdS/dS
Comparing this with the metric in eq. (98), the only difference between the induced metrics for the dS and AdS cases is the replacement of the conformal factor $\sinh^{2}w$ with $\sin^{2}w$. As a result, the subsequent calculations are identical to those in the AdS case, except for replacing $\sinh^{2}w_{c}$ with $\sin^{2}w_{c}$. Consequently, the effective radius is given by $r_{\rm eff}=\ell_{\rm dS}\sin w_{c}\cosh t_{0}\sin\phi_{a}$. For dS/dS, the entanglement entropy expression for a general spatial interval is given by eq. (50). In this case, the effective radius must be less than $\ell_{\rm dS}$, which implies
@@
Finally, we can obtain the entanglement entropy of subregion $A$ as
$S_{A}=\frac{c_{\rm dS}}{3}\arctan\left(\frac{1}{\sqrt{\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\rm eff}^{2}}-1}}\right)=\frac{c_{\rm dS}}{3}\arctan\left(\frac{1}{\sqrt{\frac{1}{(\sin w_{c}\cosh t_{0}\sin\phi_{a})^{2}}-1}}\right)\,.$ (112)
For the specific case $t=0$, we can derive the relation
$\frac{1}{(\sin w_{c}\sin\phi_{a})^{2}}-1=\cot^{2}\phi_{a}+\frac{\cot^{2}w_{c}}{\sin^{2}\phi_{a}}\,.$ (113)
Substituting this into (112), the entanglement entropy at $t_{0}=0$ becomes
$S_{A}=\frac{c_{\rm dS}}{3}\mathrm{arccot}\left(\sqrt{\cot^{2}\phi_{a}+\frac{\cot^{2}w_{c}}{\sin^{2}\phi_{a}}}\right)\,,$ (114)
@@
which is consistent with the holographic calculation of the area of the RT surface in ref. Geng:2019ruz . For the case where the subsystem is defined as $\phi\in(-\frac{\pi}{2},\frac{\pi}{2})$, the entanglement entropy is given by $S_{A}=\frac{c_{\rm dS}}{3}w_{c}$, which agrees with the result in ref. Gorbenko:2018oov .
@@
The next task is to determine whether the system’s entanglement entropy satisfies the strong subadditivity inequality. To achieve this, we compute the second-order derivative of the entanglement entropy with respect to the subsystem scale $\Delta\phi$, where $\Delta\phi=2\phi_{a}$. The result is
$\partial^{2}_{\Delta\phi}S_{A}(\Delta\phi)=-\frac{c_{\rm dS}}{12\sin^{2}\frac{\Delta\phi}{2}}\frac{\frac{1}{\sin^{2}w_{c}\cosh^{2}t_{0}}-1}{\left(\frac{1}{\sin^{2}\frac{\Delta\phi}{2}\sin^{2}w_{c}\cosh^{2}t_{0}}-1\right)^{3/2}}.$ (115)
We find that when $\sin w_{c}\cosh t_{0}>1$, the entanglement entropy always violates the strong subadditivity inequality. However, when $\sin w_{c}\cosh t_{0}<1$, for any $\Delta\phi$, the condition $r_{\rm eff}<\ell_{\rm dS}$ is always satisfied, and the entanglement entropy satisfies the strong subadditivity inequality, as illustrated in figure 12(a) and figure 12(b).
(a) $w_{c}=\frac{\pi}{2}$.
(b) $t_{0}=0$.
Figure 12: The second-order derivative of the entanglement entropy. 12(a) we take $w_{c}=\frac{\pi}{2}$, with $r_{\text{eff}}=\ell_{\rm dS}\cosh t_{0}\sin\frac{\Delta\phi}{2}$. In this case, we need to always ensure that $\cosh t_{0}\sin\frac{\Delta\phi}{2}<1$. 12(b) we take $t_{0}=0$, with $r_{\text{eff}}=\ell_{\rm dS}\sin w_{c}\sin\frac{\Delta\phi}{2}<\ell_{\rm dS}$.
For the two cases described above, the boundary positions are illustrated in figure. 13. The orange region corresponds to $\sin w_{c}\cosh t_{0}>1$, where the dual field theory is nonlocal. The blue region corresponds to $\sin w_{c}\cosh t_{0}<1$, where the dual field theory satisfies the strong subadditivity inequality. However, it can be proven that in this case, the entanglement entropy does not satisfy the boosted strong subadditivity inequality. In the dS2 background, the infinitesimal form of the boosted strong subadditivity inequality is given by Geng:2019ruz
$\text{BSSA}(\Delta\phi)=\partial^{2}_{\Delta\phi}S_{A}(\Delta\phi)+\frac{1}{\sin\Delta\phi}\partial_{\Delta\phi}S_{A}(\Delta\phi)\leq 0.$ (116)

### v3

@@ 4.1 The generalization of the replica method
If $\omega(T,X)$ takes distinct constant values at the two endpoints, denoted $\omega_{a}$ and $\omega_{-a}$, respectively, the contributions to the integral (113) from these endpoints correspond to spheres with radii $\omega_{a}X_{a}$ and $\omega_{-a}X_{a}$. These integrals are therefore replaced by the partition function on spheres with radii $\omega_{a}X_{a}$ and $\omega_{-a}X_{a}$, respectively.
@@ 4.2 System with finite size
In the preceding subsection, we reviewed the CHM mapping method for computing the entanglement entropy of $T\bar{T}$-deformed CFTs within spatial intervals on a two-dimensional plane. Specifically, we calculated the entanglement entropy of a subsystem in dS Poincaré spacetime. We now extend this methodology to two-dimensional cylindrical and spherical manifolds to investigate whether the $T\bar{T}$ deformation induces corrections to the entanglement entropy in finite-size systems.
@@ 4.2.1 AdS case
$S_{A}=$ $\frac{c_{\rm AdS}}{3}\tanh^{-1}\left(\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r_{\rm eff}^{2}}}}\right)$
@@
After the $T\bar{T}$ deformation, its holographic boundary is a two-dimensional sphere determined by $t=t_{c}$. The subregion $A$ is defined by the arc $\theta=\pi/2$, $\phi\in(-\phi_{a},\phi_{a})$ on this two-dimensional sphere. As done previously, we begin by mapping the sphere to the complex plane using the transformation
$\theta=2\tan^{-1}\sqrt{T^{2}+X^{2}}\,,\quad\phi=\frac{1}{2i}\ln\frac{T+iX}{T-iX}\,.$ (64)
@@
$ds^{2}$ $=\frac{\ell_{\rm dS}^{2}\cosh^{2}t_{c}\sin^{2}\phi_{a}}{(1+\cos(\Phi-\phi_{a})\cos\phi_{a}\sin\Theta)^{2}}(d\Theta^{2}+\sin^{2}\Theta d\Phi^{2})$
@@
$\Omega(\Theta,\Phi)=\frac{1}{1+\cos(\Phi-\phi_{a})\cos\phi_{a}\sin\Theta}\,.$ (68)
@@
$S_{A}$ $=-\frac{ic_{\rm dS}}{3}\coth^{-1}\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\rm eff}^{2}}}}\right)+\frac{\pi c_{\rm dS}}{6}$
@@
This result is consistent with the area of the RT surface calculated in Appendix B.2, and also matches the findings of ref. Chen:2023eic when $\phi_{a}=\frac{\pi}{2}$.
@@ 4.3 System with finite temperature
$S_{A}=\frac{c_{\rm AdS}}{6}\left[\tanh^{-1}\left(\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r_{{\rm eff},\Theta=0}^{2}}}}\right)+\tanh^{-1}\left(\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r_{{\rm eff},\Theta=\pi}^{2}}}}\right)\right]\,.$ (80)
@@
where we have replaced $R_{0}$ with $\frac{2\pi\ell_{\rm AdS}^{2}}{\beta}$. In the high-temperature limit $\beta\ll 1$, the last term is negligible compared to the second term [^5] This result is consistent with the findings in refs. Chen:2018eqk ; Apolo:2023ckr by identifying $R_{c}^{2}=\frac{6\ell_{\rm AdS}^{2}}{\mu\pi c_{\rm AdS}}$..
5 dS/dS holography and half-dS holography
In this section, we apply the replica method to dS/dS holography and half-dS holography. In contrast to the previously discussed examples, here, all boundary-induced metrics are dS metrics, and the dual field theory resides on dS spacetime. Previous studies on dS/dS holography and half-dS holography Kawamoto:2023nki ; Geng:2019ruz suggest that the dual field theory exhibits non-locality. In this section, we will provide a more detailed discussion of entanglement entropy and non-locality, building upon these prior investigations.
Typically, the locality of a field theory is examined through strong subadditivity (SSA) and boosted strong subadditivity (BSSA). In Appendix C, we provide a concise review of the derivation of SSA and BSSA in flat spacetime. However, since the dual field theory in dS/dS holography and half-dS holography resides on dS space, we need to account for the effects of curved spacetime. The causal structure of dS spacetime exhibits a distinct geometry compared to the Minkowski causal diamond, owing to the metric’s inherent time evolution, as illustrated in figure 12.
Figure 12: The causal structure of dS spacetime. The interval size we choose is $\Delta\phi=\frac{98\pi}{100}$.
In dS spacetime, since the SSA does not consider boost transformations, the intervals $A$ and $B$ remain on the $t=0$ slice. Consequently, the SSA is unaffected by the curvature of spacetime. Thus, the SSA in dS space is
$S^{\prime\prime}(\Delta\phi)\leq 0.$ (82)
When deriving the BSSA, we must consider a small boost transformation of intervals $A$ and $B$. In figure 12, we represent interval $A$ with red segments, interval $B$ with blue segments, $A\cap B$ with orange segments, and $A\cup B$ with green segments. Considering that the entanglement entropy is a function of the interval size, we use the geodesic distance between two points as the interval length. We set the length of interval $A\cup B$ as $L_{A\cup B}=\Delta\phi$, and consider a small boost transformation that shrinks $\delta\phi$ in the $\phi$ direction. Then, the length of interval $A\cap B$ is
$L_{A\cap B}=\cos^{-1}\left(1-2\sin^{2}\left(\frac{\Delta\phi}{2}-\delta\phi\right)\cosh\left(2\tanh^{-1}\left(\tan\frac{\delta\phi}{2}\right)\right)\right)\,.$ (83)
The lengths of intervals $A$ and $B$ are
$L_{A}=L_{B}=\cos^{-1}\left(\cos(\Delta\phi-\delta\phi)\cosh\left(2\tanh^{-1}\left(\tan\frac{\delta\phi}{2}\right)\right)\right)\,.$ (84)
Taking the limit $\delta\phi\to 0$ and expanding formula (139) in a series, we obtain the BSSA in dS space:
$S^{\prime\prime}(\Delta\phi)+\frac{1}{\sin\Delta\phi}S^{\prime}(\Delta\phi)\leq 0\,.$ (85)
It is important to note that the preceding discussion is applicable only to the constant time slice at $t=0$. For cases where $t\neq 0$, we will directly use the entanglement entropy results obtained in the text to verify whether the SSA and BSSA are satisfied.
5.1 The dS/dS holography
@@
In the Euclidean signature, after rescaling the time coordinate, the boundary metric can be written as
$ds^{2}=\ell_{\rm dS}^{2}\frac{\sin^{2}w_{c}}{\cosh^{2}\tilde{t}}\left(d\tilde{t}^{2}+d\phi^{2}\right)\,,$ (88)
where $\tilde{t}$ is related to $t$ by the relation $\tilde{t}=\ln\frac{\cot\frac{t}{2}+1}{\cot\frac{t}{2}-1}$. This metric is conformally equivalent to the cylinder. As we have done in the last section, we can map it to the complex plane through the following transformations
$\tilde{t}\rightarrow\frac{1}{2}\ln(T^{2}+X^{2})\,,\quad\phi\rightarrow\frac{1}{2i}\ln\left(\frac{T+iX}{T-iX}\right)\,.$ (89)
and the resulting conformal flat metric is
$ds^{2}=\ell_{\rm dS}^{2}\sin^{2}w_{c}\frac{4\left(dT^{2}+dX^{2}\right)}{\left(1+T^{2}+X^{2}\right)^{2}}\,.$ (90)
The two endpoints of the subregion $A$ in the $(T,X)$ coordinates are
$(T_{1},X_{1})=(e^{\tilde{t}_{0}}\cos\phi_{a},-e^{\tilde{t}_{0}}\sin\phi_{a}),\quad(T_{2},X_{2})=(e^{\tilde{t}_{0}}\cos\phi_{a},e^{\tilde{t}_{0}}\sin\phi_{a})\,.$ (91)
Using the CHM map,
$ds^{2}=\frac{4e^{2\tilde{t}_{0}}\ell_{\rm dS}^{2}\sin^{2}w_{c}\sin^{2}\phi_{a}\left(d\Theta^{2}+\sin^{2}\Theta d\Phi^{2}\right)}{\left(1+\sin\Theta\cos\Phi+e^{2\tilde{t}_{0}}\sin\Theta\cos(\Phi-2\phi_{a})+e^{2\tilde{t}_{0}}\right)^{2}}\,.$ (93)
It is easy to verify that the conformal factor at the endpoints $\Theta=0,\pi$ is the same constant, which can be absorbed into $r_{\rm eff}$. Therefore, the effective radius is $r_{\text{eff}}=\ell_{\rm dS}\sin w_{c}\cos t_{0}\sin\phi_{a}$. Finally, in Lorentz signature, the effective radius is given by
$r_{\text{eff}}=\ell_{\rm dS}\sin w_{c}\cosh t_{0}\sin\phi_{a}\,.$ (94)
As mentioned in Section 3.2, we require that the effective spherical radius be less than $\ell_{\rm dS}$, i.e.,
@@
In fact, this ensures that the geodesic length connecting the endpoints of the selected subregion in dS spacetime is real Geng:2020kxh ; Dong:2018cuv . Finally, we obtain the entanglement entropy of subregion $A$ as
$S_{A}=\frac{c_{\rm dS}}{3}\tan^{-1}\left(\frac{1}{\sqrt{\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\rm eff}^{2}}-1}}\right)=\frac{c_{\rm dS}}{3}\tan^{-1}\left(\frac{1}{\sqrt{\frac{1}{(\sin w_{c}\cosh t_{0}\sin\phi_{a})^{2}}-1}}\right)\,.$ (96)
For the specific case $t=0$, the entanglement entropy at $t_{0}=0$ becomes
$S_{A}=\frac{c_{\rm dS}}{3}\cot^{-1}\left(\sqrt{\cot^{2}\phi_{a}+\frac{\cot^{2}w_{c}}{\sin^{2}\phi_{a}}}\right)\,,$ (97)
@@
which is consistent with the holographic calculation of the area of the RT surface in ref. Geng:2019ruz . For the case where the subsystem is defined as $\phi\in(-\frac{\pi}{2},\frac{\pi}{2})$, the entanglement entropy is given by $S_{A}=\frac{c_{\rm dS}}{3}w_{c}$, which agrees with the result in ref. Gorbenko:2018oov . The subsequent task is to determine whether the system’s entanglement entropy satisfies both SSA and BSSA. We will discuss the two cases of $t_{0}=0$ and $t_{0}\neq 0$ separately.
5.1.1 The case of $t_{0}=0$
@@
The second-order derivative of the entanglement entropy with respect to the subsystem scale $\Delta\phi$ ($\Delta\phi=2\phi_{a}$) is
$\partial^{2}_{\Delta\phi}S_{A}(\Delta\phi)=-\frac{c_{dS}}{12\sin^{2}\frac{\Delta\phi}{2}}\frac{\frac{1}{\sin^{2}w_{c}}-1}{\left(\frac{1}{\sin^{2}\frac{\Delta\phi}{2}\sin^{2}w_{c}}-1\right)^{3/2}}<0.$ (98)
It can be observed that the SSA is manifestly satisfied, as illustrated in figure 13(a). However, the BSSA is violated in this case. Substituting Eq. (97) into the left-hand side of the BSSA (85), we obtain
$\text{BSSA}(\Delta\phi)=\frac{c_{dS}}{48}\frac{\sin^{2}\Delta\phi}{\sin^{6}\frac{\Delta\phi}{2}\sin^{2}w_{c}\left(\frac{1}{\sin^{2}\frac{\Delta\phi}{2}\sin^{2}w_{c}}-1\right)^{\frac{3}{2}}}>0,$ (99)
where $\text{BSSA}(\Delta\phi)$ represents $S^{\prime\prime}(\Delta\phi)+\frac{1}{\sin\Delta\phi}S^{\prime}(\Delta\phi)$. This indicates that the dual field theory continues to exhibit non-locality. Notably, at the ultraviolet boundary where $w_{c}=\frac{\pi}{2}$, there is a significant violation of the BSSA [^6] In the $T\bar{T}$-deformed AdS/CFT, the BSSA violation also occurs due to the influence of the $T\bar{T}$ deformation Aharony:2023dod . However, as the finite truncation $r_{c}$ approaches the boundary, the BSSA tends to saturate, which contrasts with the behavior in the dS spacetime background., as shown by the purple line in figure 13(b).
(a) The SSA is satisfied.
(b) The BSSA is violated.
Figure 13: The SSA and BSSA in dS/dS holography at $t_{0}=0$. 13(a) The SSA is satisfied at $t_{0}=0$. When $w_{c}=\frac{\pi}{2}$, the SSA is saturated. 13(b) The BSSA is violated at $t_{0}=0$. When $w_{c}=\frac{\pi}{2}$, the violation of the BSSA inequality is obvious.
5.1.2 The case of $t_{0}\neq 0$
In the case of $t_{0}=0$, $\sin w_{c}\cosh t_{0}<1$, but for $t_{0}\neq 0$, the value of $\sin w_{c}\cosh t_{0}$ can be either greater than 1 or less than 1. Since the effective spherical radius $r_{\text{eff}}=\ell_{dS}\sin w_{c}\cosh t_{0}\sin\frac{\Delta\phi}{2}$ is required to be less than $\ell_{dS}$, we must consider both cases, $\sin w_{c}\cosh t_{0}<1$ and $\sin w_{c}\cosh t_{0}>1$, separately.
(a) The SSA is satisfied.
(b) The BSSA is violated.
Figure 14: The SSA and BSSA in dS/dS holography at $\sin w_{c}\cosh t_{0}<1$. 14(a) The SSA is satisfied at $\sin w_{c}\cosh t_{0}<1$. The entanglement entropy, as a function of the subinterval scale $\Delta\phi$, is a concave function, ensuring that subadditivity is always satisfied. 14(b) The BSSA is violated at $\sin w_{c}\cosh t_{0}<1$. To maintain $\sin w_{c}\cosh t_{0}<1$, we set $t_{0}$ to 0.01, 0.03, 0.05, 0.07, and 0.09, respectively, and take $\delta\phi=0.01\Delta\phi$.
When $\sin w_{c}\cosh t_{0}<1$, the condition $r_{\text{eff}}<\ell_{dS}$ is always satisfied for any $\Delta\phi$. Since $t\neq 0$, we directly use the entanglement entropy to verify whether SSA and BSSA are satisfied. As shown in figure 14(a), when $\sin w_{c}\cosh t_{0}<1$, the entanglement entropy is a concave function, and the concavity of the function indicates that $\partial^{2}_{\Delta\phi}S(\Delta\phi,w_{c},t_{0})<0$, confirming that the SSA is satisfied. However, upon considering the boost transformation, we find that the BSSA is consistently violated, as shown in Figure 14(b).
(a) The SSA is violated.
(b) The BSSA is violated.
Figure 15: The SSA and BSSA in dS/dS holography at $\sin w_{c}\cosh t_{0}>1$. 15(a) The SSA is violated at $\sin w_{c}\cosh t_{0}>1$. The entanglement entropy is a convex function, ensuring that subadditivity is always violated. 15(b) The BSSA is violated at $\sin w_{c}\cosh t_{0}>1$. We set $t_{0}$ to 1.2, 1.4, 1.6, 1.8, and 2.0, respectively, and take $\delta\phi=0.1\Delta\phi$.
When $\sin w_{c}\cosh t_{0}>1$, to ensure that the effective spherical radius is less than $\ell_{\rm dS}$, the value of $\Delta\phi$ has a maximum value $\Delta\phi_{\text{max}}$ such that $\sin w_{c}\cosh t_{0}\sin\frac{\Delta\phi_{\text{max}}}{2}=1$. As shown in figure 15(a), in the interval $[0,\Delta\phi_{\text{max}}]$, the entanglement entropy as a function of $\Delta\phi$ is consistently a convex function, which ensures that the subadditivity is always violated. Similarly, considering the boost transformation, it can be observed that the BSSA is still violated when $\sin w_{c}\cosh t_{0}$ is greater than 1, as shown in figure 15(b).